# MultiConIR: Towards multi-condition Information Retrieval

### v2

@@ Appendix A Details of Models
For each model used in this paper, Table 3 provides details on model size, architecture, maximum input context length, and whether instructions is included.
@@
bge-reranker-v2-gemma (Chen et al., 2024) 2.5B Decoder Yes 1k N/A
@@
Appendix B Prompt Templates For Constructing MultiCon Dataset
Table 4, 5, and 6 present the prompts used in Steps 1 to 3 for constructing our MultiCon dataset.
[…]
@@ Appendix B Prompt Templates For Constructing MultiCon Dataset
Appendix C Details And Examples Of The MultiCon Dataset
Table 7 presents the five domains and their source data used for constructing our MultiCon dataset. Tables 9, 10, 8, 11, and 12 illustrate examples from their respective domains.
Domain Records Source Data
People 420 People Wikipedia Dat(Mahajan, 2017)
Books 482 Books Dataset (Rustamov, 2021)
Movies 500 Wikipedia Movie Plots (Robischon, 2018)
Medical Case 479 Medical Cases (HPE AI Solutions, 2023)
Legal Document 426 LexGLUE (Chalkidis et al., 2022)
Table 7: Domain-specific Source Data References. The Records column indicates the number of data entries for each domain in the MultiCon dataset.
Query 10 Positive HN 1
@@ Appendix C Details And Examples Of The MultiCon Dataset
Find a notable individual who meets these criteria: 1. Studies plasma melatonin to assess biological rhythm disorders. 2. Identified 25-hour circadian rhythms in totally blind individuals. 3. Worked at NIMH in Bethesda, Maryland before 1981. 4. Collaborated closely with Robert Sack in Oregon. 5. Researches chronobiologic sleep and mood disorders. 6. Graduated from University of Chicago in nineteen seventies with MD and PhD. 7. Had around 90 publications on PubMed by 2005. 8. Is a full professor and vice-chair of Psychiatry at OHSU. 9. Focuses on bright light exposure and melatonin treatments. 10. Studies disorders like winter depression, jet lag, and shift work maladaptation. Relying on a very precise assay for plasma melatonin, a hormone that has a clearly defined 24-hour pattern of secretion, biological rhythm disorders can be assessed and their treatment can be monitored. Totally blind individuals have 25-hour circadian rhythms, drifting an hour later each day unless they take a melatonin capsule at a certain time every day. Prior to moving to Oregon in 1981, Lewy was at the National Institute of Mental Health (NIMH) in Bethesda, Maryland, working with senior colleague Thomas Wehr. In Oregon, he has worked closely with Robert L. Sack. He describes his research as follows: ’My laboratory studies chronobiologic sleep and mood disorders.’ Alfred J. Lewy, aka Sandy Lewy, graduated from University of Chicago in 1973 after studying psychiatry, pharmacology, and ophthalmology. As of December 2005, he had 94 publications available on PubMed. He is a full professor and vice-chair of the Department of Psychiatry at OHSU (Oregon Health Science University) and holds an MD and PhD. Current research is focused on developing bright light exposure and melatonin administration as treatment modalities for these disorders. These disorders include winter depression, jet lag, maladaptation to shift work, and certain types of sleep disturbances. Relying on a very precise assay for plasma melatonin, a hormone that has a clearly defined 24-hour pattern of secretion, biological rhythm disorders can be assessed and their treatment can be monitored. Totally blind individuals have 25-hour circadian rhythms, drifting an hour later each day unless they take a melatonin capsule at a certain time every day. Prior to moving to Oregon in 1981, Lewy was at the National Institute of Mental Health (NIMH) in Bethesda, Maryland, working with senior colleague Thomas Wehr. In Oregon, he has worked closely with Robert L. Sack. He describes his research as follows: ’My laboratory studies chronobiologic sleep and mood disorders.’ Alfred J. Lewy, aka Sandy Lewy, graduated from University of Chicago in 1973 after studying psychiatry, pharmacology, and ophthalmology. As of December 2005, he had 94 publications available on PubMed. He is a full professor and vice-chair of the Department of Psychiatry at OHSU (Oregon Health Science University) and holds an MD and PhD. Current research is focused on developing bright light exposure and melatonin administration as treatment modalities for these disorders. These lifestyle changes address issues like winter depression, jet lag, shift work maladaptation, and certain types of sleep disturbances.
@@
Query 10 Positive HN 1
Find a notable individual who meets these criteria: 1. Details Bernie Madoff’s \$65B Ponzi collapse. 2. Covers the impact on national media. 3. Investigates Madoff’s history of fraud. 4. Offers deep insight into Madoff’s family. 5. Contains exclusive news and material. 6. Includes Bernie’s Little Black Book. 7. Features Madoff’s past calendars. 8. Provides details on the origins of his scam. 9. Includes cooperation from Eleanor Squillari. 10. Reveals SEC connections to Madoff. The collapse of Bernie Madoff’s Ponzi scheme led to the instant evaporation of \$65 billion of wealth. The effects of Madoff’s brazen fraud were felt most closely in New York and Palm Beach but the story was, and continues to be, front page news across the country. Brian Ross and his team of investigators shed an unyielding light onto Madoff’s scheme–how he got started, how he succeed for so long, who helped him, and who shielded him from early investigations. This is an incisive and voyeuristic look into this first family of financial crime. The Madoff Chronicles includes a vast array of news and material that readers won’t find anywhere else. Contains a reproduction of Bernie’s Little Black Book. Ross has also secured Madoff’s calendar for the past three years and other never-before-seen documents from inside the Madoff empire, straight from his desk. Read key details of how Madoff carried out his scam and the revelation that he began the fraud from almost the first day, in the 1960s. Extensive cooperation by Madoff’s personal assistant, Eleanor Squillari. Contains incriminating connections between Madoff and certain members of the SEC. The collapse of Bernie Madoff’s Ponzi scheme led to the instant evaporation of \$65 billion of wealth. The effects of Madoff’s brazen fraud were felt most closely in New York and Palm Beach but the story was, and continues to be, front page news across the country. Brian Ross and his team of investigators shed an unyielding light onto Madoff’s scheme–how he got started, how he succeed for so long, who helped him, and who shielded him from early investigations. This is an incisive and voyeuristic look into this first family of financial crime. The Madoff Chronicles includes a vast array of news and material that readers won’t find anywhere else. Contains a reproduction of Bernie’s Little Black Book. Ross has also secured Madoff’s calendar for the past three years and other never-before-seen documents from inside the Madoff empire, straight from his desk. Read key details of how Madoff carried out his scam and the revelation that he began the fraud from almost the first day, in the 1960s. Extensive cooperation by Madoff’s personal assistant, Eleanor Squillari. Contains no documented communications between Madoff and SEC members.
@@
Query 10 Positive HN 1
Find a movie that matches all conditions: 1. Originated from American. 2. Plot: Mrs. Lowe and Black were lovers. 3. Plot: Terry carelessly spends money. 4. Plot: Terry promoted to ship foreman. 5. Cast: Mary Miles Minter, Allan Forrest. 6. Plot: Julia Deep works behind exchange desk. 7. Director: Lloyd Ingraham. 8. Plot: Terry’s spending takes a toll. 9. Plot: Lottie sees Terry and Julia at park. 10. Release Year: 1918. Origin/Ethnicity: American
@@
Query 10 Positive HN 1
Find a case where the patient: 1. Underwent ascending aortic arch angiogram. 2. Had left common carotid artery angiogram. 3. Received right common carotid artery angiogram. 4. Undergone left subclavian artery angiogram. 5. Had right iliac angiogram with runoff. 6. Performed bilateral cerebral angiograms. 7. Experienced TIA and moderate carotid stenosis. 8. Had 400 ml blood loss. 9. Provided informed consent for the procedure. 10. Received a 6-French sheath in the right femoral artery. PROCEDURE PERFORMED:
@@
Appendix D Complete Results
D.1 Complete Results Of Task1
Table 13, 14, 15, 16, and 17 present the complete results of Task 1 for the five datasets: People, Books, Movies, Medical Case, and Legal Document.
Spare Retriever
BM25 23.10 25.24 22.62 26.43 26.19 23.57 24.05 22.14 22.38 23.81 -0.71
Dense Retriever
jina-embeddings-v3 88.81 78.33 78.33 76.90 72.14 68.57 67.38 66.19 67.62 65.95 22.86
gte-large-en-v1.5 72.14 76.19 73.10 68.57 69.29 61.43 66.19 62.38 69.05 63.81 8.33
NV-Embed-v2 82.14 89.52 85.48 85.48 87.38 84.05 87.38 84.05 84.05 86.67 -4.52
bge-en-icl 85.00 85.95 86.43 85.00 83.81 84.29 86.43 87.14 85.48 83.81 1.19
gte-Qwen2-7B-instruct 58.10 70.95 69.05 70.71 71.90 71.43 71.43 69.76 72.62 69.76 -11.67
gte-Qwen2-1.5B-instruct 61.90 74.76 70.24 71.19 73.57 70.48 73.81 66.90 65.71 65.00 -3.10
e5-mistral-7b-instruct 76.43 77.86 73.33 75.48 68.10 71.43 67.62 68.81 69.05 73.57 2.86
GritLM-7B 79.05 84.52 80.71 81.67 81.67 77.86 83.10 82.62 81.43 82.62 -3.57
LLM2Vec 79.05 74.76 71.90 75.71 74.05 75.95 75.00 78.81 76.19 80.00 -0.95
Reranker
bge-reranker-v2-m3 90.00 85.95 78.57 73.81 74.29 62.14 63.57 56.67 52.14 36.67 53.33
bge-reranker-v2-gemma 92.62 93.33 91.67 88.33 89.29 84.76 81.43 72.38 69.52 50.71 41.90
followIR 84.52 80.48 76.43 68.57 60.71 56.90 50.00 34.05 20.95 8.57 75.95
Table 13: Win Rate Score as the Number of Conditions in the Query Increases on the People Dataset (Task 1)
Spare Retriever
BM25 45.85 47.93 47.93 49.17 47.72 51.24 50.21 48.34 53.94 46.06 -0.21
Dense Retriever
jina-embeddings-v3 76.14 72.41 71.99 71.37 67.01 64.73 67.43 63.07 63.28 58.51 17.63
gte-large-en-v1.5 73.86 75.93 71.99 70.33 68.67 69.71 65.77 66.60 64.32 70.95 2.90
NV-Embed-v2 78.01 73.65 75.93 68.88 69.71 69.50 68.46 68.05 67.84 71.58 6.43
bge-en-icl 78.42 76.76 75.31 72.20 72.41 69.29 70.54 70.12 70.95 72.41 6.02
gte-Qwen2-7B-instruct 71.99 63.90 64.52 61.83 54.77 50.00 50.00 49.17 48.55 50.83 21.16
gte-Qwen2-1.5B-instruct 70.54 69.09 70.33 69.92 65.98 60.58 64.11 62.45 64.11 56.85 13.69
e5-mistral-7b-instruct 73.03 65.77 68.88 65.98 60.58 59.96 59.34 60.58 60.37 68.67 4.36
GritLM-7B 80.91 74.69 81.33 73.86 75.73 73.86 79.25 77.80 78.01 80.50 0.41
LLM2Vec 80.08 70.12 72.41 70.12 68.05 66.80 64.32 65.35 65.15 64.52 15.56
Reranker
bge-reranker-v2-m3 83.20 79.05 76.35 71.99 72.20 72.82 74.48 68.67 60.58 51.66 31.54
bge-reranker-v2-gemma 87.76 84.44 82.78 81.74 80.08 75.31 74.48 72.82 70.75 70.54 17.22
followIR 85.89 78.22 77.59 72.41 68.67 68.46 61.41 60.37 53.11 47.30 38.59
Table 14: Win Rate Score as the Number of Conditions in the Query Increases on the Books Dataset (Task 1)
Spare Retriever
BM25 28.60 34.60 33.60 37.20 34.80 36.80 37.00 38.40 37.60 39.00 -10.40
Dense Retriever
jina-embeddings-v3 77.60 77.00 75.20 75.40 70.80 72.00 69.40 68.40 71.80 67.20 10.40
gte-large-en-v1.5 78.40 81.00 76.60 75.60 73.80 71.00 71.00 69.00 67.20 59.40 19.00
NV-Embed-v2 81.00 81.80 80.40 78.60 80.00 76.00 77.40 76.60 72.20 67.00 14.00
bge-en-icl 84.40 83.20 80.60 79.20 78.80 77.00 76.00 75.80 70.40 66.60 17.80
gte-Qwen2-7B-instruct 74.60 80.00 74.60 72.80 72.40 72.00 68.80 75.00 67.60 67.40 7.20
gte-Qwen2-1.5B-instruct 80.40 83.00 76.00 74.80 73.00 73.80 71.00 74.60 72.20 68.40 12.00
e5-mistral-7b-instruct 73.40 71.80 66.00 64.80 70.00 61.20 65.60 62.80 63.00 55.60 17.80
GritLM-7B 85.00 83.60 80.00 80.80 79.80 72.80 76.40 72.60 71.40 73.80 11.20
LLM2Vec 87.20 89.40 86.40 84.40 81.80 81.60 79.60 77.80 75.40 74.20 13.00
Reranker
bge-reranker-v2-m3 92.80 90.80 80.20 81.60 79.20 75.60 71.00 63.40 67.00 60.40 32.40
bge-reranker-v2-gemma 92.20 91.60 86.00 89.20 83.40 79.40 77.60 78.20 70.80 61.60 30.60
followIR 92.60 86.40 83.40 88.00 81.00 81.00 83.40 81.40 82.60 85.00 7.60
Table 15: Win Rate Score as the Number of Conditions in the Query Increases on the Movies Dataset (Task 1)
Spare Retriever
BM25 25.68 28.18 31.52 32.99 37.37 36.33 35.49 39.25 37.16 40.71 -15.03
Dense Retriever
jina-embeddings-v3 68.89 64.93 65.34 65.55 55.32 59.08 56.16 63.67 60.54 59.29 9.60
gte-large-en-v1.5 77.24 81.84 76.83 75.57 71.82 70.56 70.35 67.01 68.48 62.42 14.82
NV-Embed-v2 77.24 73.07 74.53 73.70 65.97 62.63 67.43 68.68 67.22 63.67 13.57
bge-en-icl 80.79 78.91 76.41 77.04 68.89 68.68 72.23 71.19 69.31 67.64 13.15
gte-Qwen2-7B-instruct 71.61 73.70 73.49 72.65 66.60 67.64 66.18 64.72 60.75 51.57 20.04
gte-Qwen2-1.5B-instruct 73.90 72.86 72.03 70.77 67.43 65.55 63.47 64.09 60.13 53.86 20.04
e5-mistral-7b-instruct 69.31 67.01 67.64 64.09 57.83 55.53 56.16 51.98 50.10 48.64 20.67
GritLM-7B 79.75 82.25 76.62 78.08 77.66 75.16 75.16 77.87 78.91 78.29 1.46
LLM2Vec 84.34 81.21 79.12 83.09 80.58 79.75 78.71 79.33 80.79 78.50 5.85
Reranker
bge-reranker-v2-m3 81.21 91.02 86.64 86.43 79.33 75.37 75.78 66.39 72.65 62.00 19.21
bge-reranker-v2-gemma 89.35 92.48 90.19 92.90 91.02 85.80 85.39 82.67 80.79 70.35 19.00
followIR 92.28 90.81 89.35 89.56 86.22 84.97 81.42 86.22 76.41 71.82 20.46
Table 16: Win Rate Score as the Number of Conditions in the Query Increases on the Medical Case Dataset (Task 1)
Spare Retriever
BM25 19.72 34.04 30.05 36.85 40.85 41.78 43.66 43.90 43.19 49.77 -30.05
Dense Retriever
jina-embeddings-v3 69.01 63.62 68.31 65.96 62.68 63.85 63.38 59.86 59.86 52.58 16.43
gte-large-en-v1.5 77.70 71.36 70.42 61.03 69.95 64.32 64.32 59.86 57.75 50.23 27.46
NV-Embed-v2 84.27 83.57 77.70 71.83 75.35 70.89 65.73 60.33 58.69 51.17 33.10
bge-en-icl 88.50 78.40 73.47 70.42 68.78 65.73 65.96 61.97 52.35 49.53 38.97
gte-Qwen2-7B-instruct 77.46 72.54 68.31 64.55 60.33 56.57 54.69 52.35 46.24 41.31 36.15
gte-Qwen2-1.5B-instruct 81.46 75.12 72.54 70.19 69.72 66.90 62.21 55.16 56.10 49.30 32.16
e5-mistral-7b-instruct 83.10 71.83 65.02 66.90 61.97 59.86 49.77 51.17 46.71 44.13 38.97
GritLM-7B 85.68 76.53 73.24 67.61 67.14 67.84 64.55 62.21 62.91 64.55 21.13
LLM2Vec 84.98 71.60 67.61 64.08 57.98 58.69 55.16 52.35 46.01 37.79 47.18
Reranker
bge-reranker-v2-m3 88.50 80.99 71.36 66.43 66.43 56.10 54.46 42.25 25.59 13.62 74.88
bge-reranker-v2-gemma 93.43 88.26 82.86 72.77 72.07 69.72 60.56 55.40 43.66 27.23 66.20
followIR 61.74 62.68 54.46 54.46 53.99 48.36 36.85 17.61 9.86 4.93 56.81
Table 17: Win Rate Score as the Number of Conditions in the Query Increases on the Legal Documents Dataset (Task 1)
D.2 Complete Results Of Task 2

### v3

@@ Appendix A Details of Models
For each model used in this paper, Table 4 provides details on model size, architecture, maximum input context length, and whether instructions is included. The GPT-4o model utilized for dataset generation and ranking in this work was the "2024-07-01-preview" version.
@@
bge-reranker-v2-gemma (Chen et al., 2024) 2.5B Decoder Yes 8K N/A
@@
RankZephyr (Pradeep et al., 2023) 7B Decoder Yes 32K N/A
Zero-shot LLM for Ranking
GPT-4o (OpenAI, 2024) N/A Decoder Yes 128K N/A
@@
Appendix B Prompt Templates For Constructing Multiconir Dataset
Table 5, 6, and 7 present the prompts used in Steps 1 to 3 for constructing our Multiconir dataset.
For placeholders, {domain} $\in$ {People, Books, Movies, Medical Case, Legal Document}. {domain_features} specifies key attributes within a particular domain. In the medical case domain, {domain_features} $\in$ { patient symptoms, clinical diagnosis, drug allergies, family medical history, surgical details, postoperative outcomes, hospitalization duration, recovery status.} In the legal document domain, {domain_features} $\in$ { case type, involved parties, court ruling, legal provisions, evidence summary, defense strategy. } In the movies domain, {domain_features} $\in$ { summary, lead actors, release date, release area, genre, detailed plots. } In the books domain, {domain_features} $\in$ { author, publication year, genre, main content, detailed plots. } In the people domain, {domain_features} $\in$ { profession, nationality, notable achievements, social impact, related events. }
@@ Appendix B Prompt Templates For Constructing Multiconir Dataset
Appendix C Benchmark Quality Evaluation
To guarantee the reliability of MultiConIR, we applied a two–stage audit that examines both query validity and document–label validity. Figure 3 gives a visual outline; full numbers appear in Table 1.
Query validity.
From each of the five domains (People, Books, Movies, Medical, Legal) we randomly sampled 100 multi-condition queries, yielding a 500-item evaluation set. Ten trained graduate annotators independently rated every query for linguistic naturalness, precision of constraints, and contextual plausibility (i.e., whether logically plausible). Inter-annotator agreement reached $93.7\%$ with Fleiss’ $\kappa=0.84$, indicating near-perfect consensus that the automatically generated queries resemble genuine information needs.
Document–label validity.
  1. 1.
    LLM filtering. We applied GPT-4o to the entire corpus. For every positive document $d^{+}$ the model verified that all ten conditions in $C(d)$ were satisfied; for each hard-negative document $d^{-}_{k}$ it checked that exactly k1k\!-\!1 conditions held. Instances failing these criteria (false positives or false negatives) were discarded, reducing domain sizes to: People (420), Books (482), Movies (500), Medical (479), Legal (426).
  2. 2.
    Human spot-check. To confirm the LLM filter, we randomly drew another 100 document–query pairs per domain (500 in total). Two independent annotators judged whether the labelled number of satisfied conditions was correct; disagreements were resolved by a third adjudicator. The residual error rate was $2.4\%$, implying that the automatic filter removed the vast majority of mis-labelled items.
Appendix D Discussion: Win Rate vs. Traditional IR Metrics
Why introduce Win Rate?
MultiConIR poses multi-condition queries for which models must sense fine-grained semantic differences. We focus on two abilities: (i) discriminating the positive document from a hard negative as the number of query conditions grows (Task 1); (ii) preserving a monotonic ordering in which a document satisfying $k$ conditions outranks one satisfying k1k\!-\!1 under the same query (Task 2).
Limitations of conventional metrics.
  • Precision@1 coincides with Win Rate in Task 1 (one positive vs. one HN) but, in Task 2, observes only the top result and ignores the intended hierarchy $d^{+}\succ\mathrm{HN}_{1}\succ\cdots\succ\mathrm{HN}_{10}$.
  • NDCG@$k$ introduces graded relevance but still weights absolute rank more than pairwise consistency, thus blurring step-wise violations of the monotonic order.
  • Recall proved even less informative in early pilot runs dominated by easy negatives: high recall was achievable without respecting the semantic precision that MultiConIR is designed to test.
How Win Rate fills the gap.
Win Rate computes the proportion of pairwise comparisons in which a document that fulfils more conditions is ranked above one that fulfils fewer. Hence, it matches Precision@1 in the degenerate Task 1 case, yet remains sensitive to every local inversion in the graded Task 2 ladder, offering a sharper lens on a model’s ability to capture incremental semantic constraints.
Appendix E Examples Of The MultiConIR Dataset
Tables 9, 10, 8, 11, and 12 illustrate examples from their respective domains.
Query 3 Positive HN 1
@@ Appendix E Examples Of The MultiConIR Dataset
Find a notable individual who meets these criteria: 1. Studies plasma melatonin to assess biological rhythm disorders. 2. Identified 25-hour circadian rhythms in totally blind individuals. 3. Worked at NIMH in Bethesda, Maryland before 1981. Relying on a very precise assay for plasma melatonin, a hormone that has a clearly defined 24-hour pattern of secretion, biological rhythm disorders can be assessed and their treatment can be monitored. Totally blind individuals have 25-hour circadian rhythms, drifting an hour later each day unless they take a melatonin capsule at a certain time every day. Prior to moving to Oregon in 1981, Lewy was at the National Institute of Mental Health (NIMH) in Bethesda, Maryland, working with senior colleague Thomas Wehr. In Oregon, he has worked closely with Robert L. Sack. He describes his research as follows: ’My laboratory studies chronobiologic sleep and mood disorders.’ Alfred J. Lewy, aka Sandy Lewy, graduated from University of Chicago in 1973 after studying psychiatry, pharmacology, and ophthalmology. As of December 2005, he had 94 publications available on PubMed. He is a full professor and vice-chair of the Department of Psychiatry at OHSU (Oregon Health Science University) and holds an MD and PhD. Current research is focused on developing bright light exposure and melatonin administration as treatment modalities for these disorders. These disorders include winter depression, jet lag, maladaptation to shift work, and certain types of sleep disturbances. Relying on a very precise assay for plasma melatonin, a hormone that has a clearly defined 24-hour pattern of secretion, biological rhythm disorders can be assessed and their treatment can be monitored. Totally blind individuals have 25-hour circadian rhythms, drifting an hour later each day unless they take a melatonin capsule at a certain time every day. After to moving to Oregon in 1981, Lewy was at the National Institute of Mental Health (NIMH) in Bethesda, Maryland, working with senior colleague Thomas Wehr. In Oregon, he has worked closely with Robert L. Sack. He describes his research as follows: ’My laboratory studies chronobiologic sleep and mood disorders.’ Alfred J. Lewy, aka Sandy Lewy, graduated from University of Chicago in 1973 after studying psychiatry, pharmacology, and ophthalmology. As of December 2005, he had 94 publications available on PubMed. He is a full professor and vice-chair of the Department of Psychiatry at OHSU (Oregon Health Science University) and holds an MD and PhD. Current research is focused on developing bright light exposure and melatonin administration as treatment modalities for these disorders. These disorders include winter depression, jet lag, maladaptation to shift work, and certain types of sleep disturbances.
@@
Query 5 Positive HN 1
Find a notable individual who meets these criteria: 1. Details Bernie Madoff’s \$65B Ponzi collapse. 2. Covers the impact on national media. 3. Investigates Madoff’s history of fraud. 4. Offers deep insight into Madoff’s family. 5. Contains exclusive news and material. The collapse of Bernie Madoff’s Ponzi scheme led to the instant evaporation of \$65 billion of wealth. The effects of Madoff’s brazen fraud were felt most closely in New York and Palm Beach but the story was, and continues to be, front page news across the country. Brian Ross and his team of investigators shed an unyielding light onto Madoff’s scheme–how he got started, how he succeed for so long, who helped him, and who shielded him from early investigations. This is an incisive and voyeuristic look into this first family of financial crime. The Madoff Chronicles includes a vast array of news and material that readers won’t find anywhere else. Contains a reproduction of Bernie’s Little Black Book. Ross has also secured Madoff’s calendar for the past three years and other never-before-seen documents from inside the Madoff empire, straight from his desk. Read key details of how Madoff carried out his scam and the revelation that he began the fraud from almost the first day, in the 1960s. Extensive cooperation by Madoff’s personal assistant, Eleanor Squillari. Contains incriminating connections between Madoff and certain members of the SEC. The collapse of Bernie Madoff’s Ponzi scheme led to the instant evaporation of \$65 billion of wealth. The effects of Madoff’s brazen fraud were felt most closely in New York and Palm Beach but the story was, and continues to be, front page news across the country. Brian Ross and his team of investigators shed an unyielding light onto Madoff’s scheme–how he got started, how he succeed for so long, who helped him, and who shielded him from early investigations. This is an incisive and voyeuristic look into this first family of financial crime. The Madoff Chronicles includes a vast array of news and material. Contains a reproduction of Bernie’s Little Black Book. Ross has also secured Madoff’s calendar for the past three years and other never-before-seen documents from inside the Madoff empire, straight from his desk. Read key details of how Madoff carried out his scam and the revelation that he began the fraud from almost the first day, in the 1960s. Extensive cooperation by Madoff’s personal assistant, Eleanor Squillari. Contains incriminating connections between Madoff and certain members of the SEC.
@@
Query 8 Positive HN 1
Find a movie that matches all conditions: 1. Originated from American. 2. Plot: Mrs. Lowe and Black were lovers. 3. Plot: Terry carelessly spends money. 4. Plot: Terry promoted to ship foreman. 5. Cast: Mary Miles Minter, Allan Forrest. 6. Plot: Julia Deep works behind exchange desk. 7. Director: Lloyd Ingraham. 8. Plot: Terry’s spending takes a toll. Origin/Ethnicity: American
@@
Query 7 Positive HN 1
Find a case where the patient: 1. Underwent ascending aortic arch angiogram. 2. Had left common carotid artery angiogram. 3. Received right common carotid artery angiogram. 4. Undergone left subclavian artery angiogram. 5. Had right iliac angiogram with runoff. 6. Performed bilateral cerebral angiograms. 7. Experienced TIA and moderate carotid stenosis. PROCEDURE PERFORMED:
@@
INDICATIONS FOR PROCEDURE: TIA, aortic stenosis, postoperative procedure. Severe carotid artery stenosis.
@@
Appendix F Complete Results
F.1 Complete Results Of Task 2